# A Variational Framework for Residual-Based Adaptivity in Neural PDE Solvers and Operator Learning

### v1

@@ 1 Introduction
Scientific machine learning (SciML) has emerged as a powerful alternative to traditional numerical methods for solving partial differential equations (PDEs). Here, we consider two of the main approaches in SciML. The first, which includes Physics-Informed Neural Networks (PINNs) [1] and their variants [2], focuses on function approximation, where a representation model is trained to satisfy the governing equations of a specific problem. The second is operator learning [3, 4], where a model learns the underlying solution operator itself, allowing it to generate solutions for new boundary conditions, source terms, or parameters almost instantaneously.
@@
Our unified framework provides a principled origin for heuristics like RBA and RAD, showing how different potential functions correspond to different adaptive schemes. We extend these methods to operator learning with a hybrid strategy employing importance sampling over the function space and importance weighting over the spatial domain, which can be seamlessly integrated into architectures like FNO and DeepONet. The framework yields a twofold benefit: it lowers the discretization error by reducing the variance of the loss estimator, and it improves learning dynamics by enhancing the signal-to-noise ratio of the gradients, leading to faster convergence.Finally, we demonstrate the efficacy of vRBA across a range of challenging PINN and operator learning tasks. Our empirical results show that using vRBA is critical to achieve lower errors, providing significant improvements even when paired with state-of-the-art second-order optimizers [15] or specialized architectures like TC-UNet.
@@ 2 Problem Setup
For a domain $\Omega\subset\mathbb{R}^{d}$, define the residual $r:\Omega\to\mathbb{R}_{+}$ to be a bounded function describing the error at each spatial point $x\in\Omega$. For example, in supervised learning tasks such as function approximation, one seeks to approximate some function $\bar{u}:\Omega\to\mathbb{R}$ with a parameterized $u(x;\theta)$, for parameters $\theta$ in some parameter space $\mathcal{T}$. Then, the residual is defined to be the difference:
@@
On the other hand, PINNs, the residual is given by an appropriate differential operator $\mathcal{F}$ applied onto the parameterized function:
@@
There are many possible choices of loss functions. In the ideal scenario, one would like to find the parameter that minimizes the residual uniformly over all points in the spatial domain, i.e., we seek the solution to the optimization problem:
@@
In other words, we wish to minimize the $L^{\infty}(\Omega)$-norm of the residual. Alternatively, we can (superfluously) write (P1) as an optimization over the space of probability measures:
@@
(P1) is rarely used in practice due to its instability and non-differentiability. Moreover, the optimization landscape induced by the parameterization is extremely non-convex and identically-zero residual is impossible for most parameterization schemes, e.g., neural networks, despite the availability of asymptotic approximation theorems [40, 41, 42]. Consequently, the popular alternative is to solve a “weaker” problem by minimizing:
@@
which can be easily approximated via Monte Carlo integration, that is, let $(X_{i})_{i=1}^{N}$ be independent samples from $p$, we define the discrete loss
@@ 2.1 Reformulating loss function for adaptive training
One family of methods involves sampling the quadrature points according to a suitable residual-based distribution [19, 20, 27, 28, 43, 44, 45, 46, 47, 48, 49, 50, 51]. A widely adopted example is the residual-based adaptive distribution (RAD) [20], where instead of sampling the quadrature points from the base distribution $p$, one would sample from an adaptive, tilted distribution computed as:
@@
$\mathcal{L}_{W}^{(D)}(\theta)=\frac{1}{N}\sum_{i=1}^{N}[\lambda_{i}r(X_{i},\theta)]^{2}$
while numerous proposals have been made for the specific form of $\lambda_{i}$ [21, 52, 25, 53, 24, 22, 54, 55, 56, 35]. One of the most effective methods define the multipliers explicitly from the residual, as in the residual-based-attention (RBA) method: for the spatial point $X_{i}$,
@@ 2.2 Learning Dynamics
Once the objective function ($\mathcal{L}$) is calculated, the model parameters are generally optimized using a line search algorithm of the form:
@@
where $\alpha^{k}$ is the step size, and $p^{k}=-H_{k}\nabla_{\theta}\mathcal{L}(\theta^{k})$ is the update direction which depends on the gradient of the objective function and some symmetric matrix $H_{k}$ [15]. For first-order optimizers such as Adam [57], $H_{k}$ is treated as the identity matrix. Therefore, for the discrete case, the update direction induced by Equation 5 is given by:
@@
For a discrete domain $D$ of size $N=Mb$, define $\{B_{j}\}_{j=1}^{b}$ to be an equal-sized partition of $D$, that is:
@@
Letting $\hat{p}_{j}^{k}=-\nabla_{\theta}\mathcal{L}^{(B_{j})}(\theta^{k})$ be the update direction using the batch $B_{j}$, the full-batch update direction is the average of the update directions of its partitions:
@@ 2.3 Signal-to-Noise Ratio (SNR)
As shown in the preceding section, gradient disagreement is not merely an artifact of stochastic sampling; it is an inherent property of the loss landscape, observable even within a single, complete batch of data. To analyze these variations, several studies [58, 59, 60] have considered the signal-to-noise ratio (SNR). A common formulation for the SNR of a gradient vector is:
@@
To formalize this, we can decompose the update direction $\hat{p}_{j}^{k}$ from a partition $B_{j}$ relative to the true update direction from the continuous loss in Equation (4), which we define as our “signal”: $\hat{p}^{k}=-\nabla_{\theta}\mathcal{L}^{(D)}(\theta^{k})$. The partition-based direction is thus:
@@ 3 A framework for approximate uniform minimization
In this work, we propose a dual formulation of (P1) roughly of the form
@@ 3.1 A dual reformulation
To arrive at the dual form of (P1), we use the Laplace principle (A.1), which writes the maximum as a limit of integrating an increasing singular exponential functional:
@@
To arrive at the desired form, we further express the inner functional using the Gibbs variational formula (A.2):
@@
$\mathbf{H}(q|p)=\begin{dcases}\int_{\Omega}\log\frac{dq}{dp}(x)q(dx)&\text{if}~\frac{dq}{dp}~\text{exists},\\ +\infty&\text{otherwise}.\end{dcases}$ (11)
The innermost optimization admits a closed-form solution (see A.2):
@@
The Laplace principle, referred to also as Varadhan’s lemma [64, Theorem 4.3.1], has a long history in the area of statistical physics and large deviations theory. In combination with a theorem from Bryc and the Gibbs variational principle, the Laplace principle gives an alternative characterization for large deviations principles and is foundational for the development of the weak convergence approach to large deviations of Dupuis and Ellis [65, 66]; the computation used to derive the dual form is reminiscent of the techniques used in this weak convergence approach. Similar exponential functionals are known in the community as the softmax function, commonly used in machine learning, Bayesian statistics, and transformers as a smooth approximation of the maximum function. A similar family of probability distributions was recently used by Alberts and Bilionis [67] for uncertainty quantification. The inverse temperature $\beta$ plays the role of $1/\epsilon$ in this manuscript.
@@
$\begin{dcases}&q^{k+1}(dx)\leftarrow\frac{1}{\mathsf{Z}^{k}}\exp\left(\frac{r(x;\theta^{k})}{\epsilon^{k}}\right)p(dx)~~\text{where}~~\mathsf{Z}^{k}=\int_{\Omega}\exp\left(\frac{r(x;\theta^{k})}{\epsilon^{k}}\right)p(dx);\\ &\theta^{k+1}\leftarrow\mathsf{LineSearch}(\theta^{k},q^{k+1});\\ &\epsilon^{k+1}\leftarrow\mathsf{AnnealingScheme}(k,\epsilon^{k},\theta^{k+1},q^{k+1}).\end{dcases}$ (12)
There are two design choices. The oracle $\mathsf{LineSearch}$ refers to line search (8) such as gradient descent and various higher-order methods. On the other hand, $\mathsf{AnnealingScheme}$ refers to a schedule that incrementally decreases $\epsilon$. For every fixed $\theta$ and $q$, the objective function in (P1*) is maximized by taking $\epsilon\to 0$ (A.1). However, since the optimization in $\theta$ takes place incrementally, the outer-loop optimization ought to take place in accordance with the inner loop. For example, an annealing scheme used in numerical experiments is of the form
@@
We contrast the proposed algorithm with simulated annealing [68, 69, 70]—which was the source of the terminological choice and the annealing schedule [70, Theorem 1]—where given a sufficiently smooth potential $V:\Theta\to\mathbb{R}$, one finds the solution to $\min_{\Theta}V$ to be the long-time behavior ($t\to\infty$) of the solution to the stochastic differential equation
@@ 3.2 Generalization of dual formulation
$\mathbf{D}_{\Phi}(q|p)\coloneqq\begin{dcases}\int_{\Omega}\Phi\left(\frac{dq}{dp}(x)\right)p(dx)&\text{if}~\frac{dq}{dp}~\text{exists},\\ +\infty&\text{otherwise.}\end{dcases}$
@@
    is satisfied for some $p,q\in(1,\infty)$—which includes the case where $\Phi(r)=r^{p}_{c}$—then by [72, Equation 3], we have
@@ 4.1.1 Update the tilted distribution
While the generalized Gibbs variational principle does not typically yield a closed-form update rule for the distribution $q$, one can obtain the optimal tilt by choosing specific potentials $\Phi$ and annealing schedules (to be discussed in the coming subsection). Under the appropriate choices, $q$ is proportional to the derivative of the potential $\Phi^{\prime}$. For a discrete set of points, the optimal probability mass function (p.m.f.) for the next iteration, $q^{k+1}$, is given by:
@@
This choice of potential recovers the Laplace principle, which is associated with minimizing the $L^{\infty}$ norm, and there is an exact convex duality. The derivative is $\Phi^{\prime}(x)=e^{x}$, so the optimal distribution is proportional to $\exp(|r|/\epsilon)$ under no additional assumptions. This resulting distribution is given by a variation of the softmax of the scaled residuals akin to attention mechanisms popularized in the transformer architecture [73]:
@@
In this case, the derivative is $\Phi^{\prime}(x)=2x$, which is associated with minimizing the $L^{2}$ norm. In this case, the convex duality is not exact, and the appropriate annealing schedule will be chosen shortly. The discretized p.m.f. takes the form:
@@
Furthermore, we introduce an additional smoothing mechanism by interpolating between the adaptive distribution $q^{k+1}$ and the base uniform distribution $p_{u}$. The combined update rule for the importance weights is:
@@ 4.1.2 Update the model parameters
In this approach, the weights are first normalized to recover a smooth probability mass function (p.m.f.) over the discrete domain $\Omega$:
@@
This new distribution $\bar{q}$ is then used to resample a new set of training points $\{X_{i}\}_{i=1}^{N}$ from the full collocation set $\Omega$. By focusing the sampling on high-importance regions, the loss can be computed as a standard, unweighted mean squared error on this new, more challenging set of points:
$\mathcal{L}(\theta^{k+1})=\frac{1}{N}\sum_{i=1}^{N}r(X_{i},\theta^{k+1})^{2}$ (22)
@@
Alternatively, the weights can be used directly in an importance weighting scheme. In this case, the training points $\{X_{i}\}_{i=1}^{N}$ are sampled uniformly from $\Omega$. The weights $\{\lambda_{i}^{k+1}\}$ are then applied directly to the residuals within the loss calculation, creating a weighted objective:
@@ 4.1.3 Update the regularization parameter
For this choice, there is no requirements on the choice of $\epsilon$. The particular choice we implemented is:
@@
This schedule has several advantages. Using the maximum residual, $\max_{\Omega}|r|$, in the numerator provides a dynamic, problem-dependent characteristic scale for the temperature $\epsilon$. This helps to normalize the magnitude of the residuals relative to the magnitude of the solution itself. The logarithmic term in the denominator ensures a slow and stable decay, such that $\epsilon^{k}\to 0$ as $k\to\infty$, which gradually sharpens the distribution’s focus on the largest residuals. In the context of simulated annealing, logarithmic decay is sufficiently slow to guarantee global convergence [70, Theorem 1].
@@
For this case, the constraint is satisfied by choosing $\epsilon^{k}$ to be the normalizing constant, i.e., the optimal tilt is:

### v2

@@ 1 Introduction
Scientific machine learning (SciML) has emerged as a powerful alternative to traditional numerical methods for solving partial differential equations (PDEs). Here, we consider two of the main approaches in SciML. The first, which includes Physics-Informed Neural Networks (PINNs) [1] and their variants [2], focuses on function approximation, where a representation model is trained to satisfy the governing equations of a specific problem [3]. The second is operator learning [4, 5], where a model learns the underlying solution operator itself, allowing it to generate solutions for new boundary conditions, source terms, or parameters almost instantaneously.
@@
Our unified framework provides a principled origin for heuristics like RBA and RAD, showing how different potential functions correspond to different adaptive schemes. We extend these methods to operator learning with a hybrid strategy employing importance sampling over the function space and importance weighting over the spatial domain, which can be seamlessly integrated into architectures like FNO and DeepONet. The framework yields a twofold benefit: it lowers the discretization error by reducing the variance of the loss estimator, and it improves learning dynamics by enhancing the signal-to-noise ratio of the gradients, leading to faster convergence.Finally, we demonstrate the efficacy of vRBA across a range of challenging PINN and operator learning tasks. Our empirical results show that using vRBA is critical to achieve lower errors, providing significant improvements even when paired with state-of-the-art second-order optimizers [16] or specialized architectures like TC-UNet [41].
@@ 2 Problem Setup
For a domain $\Omega\subset\mathbb{R}^{d}$, define the residual $r:\Omega\to\mathbb{R}_{+}$ to be a bounded function describing the error at each spatial point $x\in\Omega$. For example, in supervised learning tasks such as function approximation, one seeks to approximate some function $\bar{u}:\Omega\to\mathbb{R}$ with a parameterized $u(x;\theta)$, for parameters $\theta$ in some parameter space $\mathcal{T}$. Then, the residual is defined as follows
@@
On the other hand, PINNs, the residual is given by an appropriate differential operator $\mathcal{F}$ applied onto the parameterized function which reads
@@
There are many possible choices of loss functions. In the ideal scenario, one would like to find the parameter that minimizes the residual uniformly over all points in the spatial domain, i.e., we seek the solution to the following optimization problem
@@
In other words, we wish to minimize the $L^{\infty}(\Omega)$-norm of the residual. Alternatively, we can (superfluously) write (P1) as an optimization over the space of probability measures
@@
(P1) is rarely used in practice due to its instability and non-differentiability. Moreover, the optimization landscape induced by the parameterization is extremely non-convex and identically-zero residual is impossible for most parameterization schemes, e.g., neural networks, despite the availability of asymptotic approximation theorems [42, 43, 44]. Consequently, the popular alternative is to solve a “weaker” problem by solving
@@
which can be easily approximated via Monte Carlo integration, that is, let $(X_{i})_{i=1}^{N}$ be independent samples from $p$, and define the discrete loss
@@ 2.1 Reformulating loss function for adaptive training
One family of methods involves sampling the quadrature points according to a suitable residual-based distribution [20, 21, 28, 29, 45, 46, 47, 48, 49, 50, 51, 52, 53]. A widely adopted example is the residual-based adaptive distribution (RAD) [21], where instead of sampling the quadrature points from the base distribution $p$, one would sample from an adaptive, tilted distribution computed as follows
@@
$\mathcal{L}_{W}^{(D)}(\theta)=\frac{1}{N}\sum_{i=1}^{N}[\lambda_{i}r(X_{i},\theta)]^{2},$
while numerous proposals have been made for the specific form of $\lambda_{i}$ [22, 54, 26, 55, 25, 23, 56, 57, 58, 36]. One of the most effective methods define the multipliers explicitly from the residual, as in the residual-based-attention (RBA) method: for the spatial point $X_{i}$, we set
@@ 2.2 Learning Dynamics
Once the objective function ($\mathcal{L}$) is calculated, the model parameters are generally optimized using a line search algorithm of the form
@@
where $\alpha^{k}$ is the step size, and $p^{k}=-H_{k}\nabla_{\theta}\mathcal{L}(\theta^{k})$ is the update direction which depends on the gradient of the objective function and some symmetric matrix $H_{k}$ [16]. For first-order optimizers such as Adam [59], $H_{k}$ is treated as the identity matrix. Therefore, for the discrete case, the update direction induced by Equation (5) is given by
@@
For a discrete domain $D$ of size $N=Mb$, define $\{B_{j}\}_{j=1}^{b}$ to be an equal-sized partition of $D$, that is
@@
Letting $\hat{p}_{j}^{k}=-\nabla_{\theta}\mathcal{L}^{(B_{j})}(\theta^{k})$ be the update direction using the batch $B_{j}$, the full-batch update direction is the average of the update directions of its partitions
@@ 2.3 Signal-to-Noise Ratio (SNR)
As shown in the preceding section, gradient disagreement is not merely an artifact of stochastic sampling; it is an inherent property of the loss landscape, observable even within a single, complete batch of data. To analyze these variations, several studies [60, 61, 62] have considered the signal-to-noise ratio (SNR). A common formulation for the SNR of a gradient vector reads
@@
To formalize this, we can decompose the update direction $\hat{p}_{j}^{k}$ from a partition $B_{j}$ relative to the true update direction from the continuous loss in Equation (4), which we define as our “signal”: $\hat{p}^{k}=-\nabla_{\theta}\mathcal{L}^{(D)}(\theta^{k})$. The partition-based direction is
@@ 3 A framework for approximate uniform minimization
In this work, we propose a dual formulation of (P1) inspired by the following problem
@@ 3.1 A dual reformulation
To arrive at the dual form of (P1), we use the Laplace principle (A.1), which writes the maximum as a limit of integrating an increasing singular exponential functional
@@
To arrive at the desired form, we further express the inner functional using the Gibbs variational formula (A.2)
@@
$\mathbf{H}(q|p)=\begin{dcases}\int_{\Omega}\log\frac{dq}{dp}(x)q(dx)&\text{if the Radon-Nikodym derivative}~\frac{dq}{dp}~\text{exists},\\ +\infty&\text{otherwise}.\end{dcases}$ (11)
The innermost optimization admits a closed-form solution (see A.2) which reads
@@
The Laplace principle, referred to also as Varadhan’s lemma [65, Theorem 4.3.1], has a long history in the area of statistical physics and large deviations theory. In combination with a theorem from Bryc and the Gibbs variational principle, the Laplace principle gives an alternative characterization for large deviations principles and is foundational for the development of the weak convergence approach to large deviations of Dupuis and Ellis [66, 67]; the computation used to derive the dual form is reminiscent of the techniques used in this weak convergence approach. Similar exponential functionals are known in the community as the softmax function, commonly used in machine learning, Bayesian statistics, and transformers as a smooth approximation of the maximum function. A similar family of probability distributions was recently used by Alberts and Bilionis [68] for uncertainty quantification. The inverse temperature $\beta$ plays the role of $1/\epsilon$ in this manuscript.
@@
$\begin{dcases}&q^{k+1}(dx)\leftarrow\frac{1}{\mathsf{Z}^{k}}\exp\left(\frac{r(x;\theta^{k})}{\epsilon^{k}}\right)p(dx)~~\text{where}~~\mathsf{Z}^{k}=\int_{\Omega}\exp\left(\frac{r(x;\theta^{k})}{\epsilon^{k}}\right)p(dx);\\ &\theta^{k+1}\leftarrow\mathsf{Minimize}\int_{\Omega}r(x,\theta^{k})q^{k+1}(dx);\\ &\epsilon^{k+1}\leftarrow\mathsf{AnnealingScheme}(k,\epsilon^{k},\theta^{k+1},q^{k+1}).\end{dcases}$ (12)
There are two design choices. The oracle $\mathsf{Minimize}$ refers to equation (8) such as gradient descent and various higher-order methods. On the other hand, $\mathsf{AnnealingScheme}$ refers to a schedule that incrementally decreases $\epsilon$. For every fixed $\theta$ and $q$, the objective function in (P1*) is maximized by taking $\epsilon\to 0$ (A.1). However, since the optimization in $\theta$ takes place incrementally, the outer-loop optimization ought to take place in accordance with the inner loop. For example, an annealing scheme used in numerical experiments is of the form
@@
We contrast the proposed algorithm with simulated annealing [69, 70, 71]—which was the source of the terminological choice and the annealing schedule [71, Theorem 1]—where given a sufficiently smooth potential $V:\Theta\to\mathbb{R}$, one finds the solution to $\min_{\Theta}V$ to be the long-time behavior ($t\to\infty$) of the solution to the stochastic differential equation
@@ 3.2 Generalization of dual formulation
$\mathbf{D}_{\Phi}(q|p)\coloneqq\begin{dcases}\int_{\Omega}\Phi\left(\frac{dq}{dp}(x)\right)p(dx)&\text{if the Radon-Nikodym derivative}~\frac{dq}{dp}~\text{exists},\\ +\infty&\text{otherwise.}\end{dcases}$
@@
    is satisfied for some $p,q\in(1,\infty)$—which includes the case where $\Phi(r)=r^{p}_{c}$—then by [73, Equation 3], we have
@@ 4.1.1 Update the tilted distribution
While the generalized Gibbs variational principle does not typically yield a closed-form update rule for the distribution $q$, one can obtain the optimal tilt by choosing specific potentials $\Phi$ and annealing schedules (to be discussed in the coming subsection). Under the appropriate choices, $q$ is proportional to the derivative of the potential $\Phi^{\prime}$. For a discrete set of points, the optimal probability mass function (p.m.f.) for the next iteration, $q^{k+1}$, is given by
@@
This choice of potential recovers the Laplace principle, which is associated with minimizing the $L^{\infty}$ norm, and there is an exact convex duality. The derivative is $\Phi^{\prime}(x)=e^{x}$, so the optimal distribution is proportional to $\exp(|r|/\epsilon)$ under no additional assumptions. This resulting distribution is given by a variation of the softmax of the scaled residuals akin to attention mechanisms popularized in the transformer architecture [74]
@@
In this case, the derivative is $\Phi^{\prime}(x)=2x$, which is associated with minimizing the $L^{2}$ norm. In this case, the convex duality is not exact, and the appropriate annealing schedule will be chosen shortly. The discretized p.m.f. takes the following form
@@
Furthermore, we introduce an additional smoothing mechanism by interpolating between the adaptive distribution $q^{k+1}$ and the base uniform distribution $p_{u}$. The combined update rule for the importance weights reads
@@ 4.1.2 Update the model parameters
In this approach, the weights are first normalized to recover a smooth probability mass function (p.m.f.) over the discrete domain $\Omega$
@@
This new distribution $\bar{q}$ is then used to resample a new set of training points $\{X_{i}\}_{i=1}^{N}$ from the full collocation set $\Omega$. By focusing the sampling on high-importance regions, the loss can be computed as a standard, unweighted mean squared error on this new, more challenging set of points
$\mathcal{L}(\theta^{k+1})=\frac{1}{N}\sum_{i=1}^{N}r(X_{i},\theta^{k+1})^{2}.$ (22)
@@
Alternatively, the weights can be used directly in an importance weighting scheme. In this case, the training points $\{X_{i}\}_{i=1}^{N}$ are sampled uniformly from $\Omega$. The weights $\{\lambda_{i}^{k+1}\}$ are then applied directly to the residuals within the loss calculation, creating a weighted objective that reads
@@ 4.1.3 Update the regularization parameter
For this choice, there is no requirements on the choice of $\epsilon$. The particular choice we implemented reads
@@
This schedule has several advantages. Using the maximum residual, $\max_{\Omega}|r|$, in the numerator provides a dynamic, problem-dependent characteristic scale for the temperature $\epsilon$. This helps to normalize the magnitude of the residuals relative to the magnitude of the solution itself. The logarithmic term in the denominator ensures a slow and stable decay, such that $\epsilon^{k}\to 0$ as $k\to\infty$, which gradually sharpens the distribution’s focus on the largest residuals. In the context of simulated annealing, logarithmic decay is sufficiently slow to guarantee global convergence [71, Theorem 1].
@@
For this case, the constraint is satisfied by choosing $\epsilon^{k}$ to be the normalizing constant, i.e., the optimal tilt is